# VideoQA-SC: Adaptive Semantic Communication for Video Question Answering

### v1

@@ Abstract
Although semantic communication (SC) has shown its potential in efficiently transmitting multi-modal data such as text, speeches and images, SC for videos has focused primarily on pixel-level reconstruction. However, these SC systems may be suboptimal for downstream intelligent tasks. Moreover, SC systems without pixel-level video reconstruction present advantages by achieving higher bandwidth efficiency and real-time performance of various intelligent tasks. The difficulty in such system design lies in the extraction of task-related compact semantic representations and their accurate delivery over noisy channels. In this paper, we propose an end-to-end SC system for video question answering (VideoQA) tasks called VideoQA-SC. Our goal is to accomplish VideoQA tasks directly based on video semantics over noisy or fading wireless channels, bypassing the need for video reconstruction at the receiver. To this end, we develop a spatiotemporal semantic encoder for effective video semantic extraction, and a learning-based bandwidth-adaptive deep joint source-channel coding (DJSCC) scheme for efficient and robust video semantic transmission. Experiments demonstrate that VideoQA-SC outperforms traditional and advanced DJSCC-based SC systems that rely on video reconstruction at the receiver under a wide range of channel conditions and bandwidth constraints. In particular, when the signal-to-noise ratio is low, VideoQA-SC can improve the answer accuracy by 5.17% while saving almost 99.5% of the bandwidth at the same time, compared with the advanced DJSCC-based SC system. Our results show the great potential of task-oriented SC system design for video applications.
@@
Keywords Semantic communication  $\cdot$ video question answering  $\cdot$ DJSCC  $\cdot$ bandwidth allocation  $\cdot$ multimodal task
@@
With the development of artificial intelligence (AI) technology, many edge machines deploy AI models to process information for intelligent tasks[1, 2], which support AI-empowered applications such as remote healthcare, autonomous driving, and the Internet of Things (IoT). Semantic communication (SC) is an emerging paradigm which aims to extract task-relevant crucial information and accomplish accurate semantic delivery, ultimately completing the intelligent tasks[3]. Thanks to the effective semantic extraction by deep neural networks, SC can achieve higher data compression ratios and faster execution of intelligent tasks compared to the traditional communication. As a result, SC is widely used in many intelligent applications that require low latency and high accuracy under limited bandwidth resources, e.g., IoT networks[4, 5], intelligently connected vehicle networks[6, 7] and smart factories [8].
@@
In a typical SC system, the transceiver is designed as a semantic codec (semantic encoder/decoder) represented by a neural network [9, 10, 11]. The semantic encoder at the transmitter needs to remove data redundancy and extract the compact semantic representation based on the structural characteristics of the source data. The semantic decoder at the receiver aims to process received semantic information to obtain results according to the specific intelligent task.
@@
For different source data modalities (speeches, text, images, etc.), appropriate neural network architectures are essential for semantic codecs to achieve efficient SC. Long Short-Term Memorys and Transformers can be utilized to model the sequential information for text[12, 13], and convolutional neural networks can be utilized to extract local information for speeches[9, 14, 15] and images[16, 10, 17]. Furthermore, deep joint source-channel coding (DJSCC) can be integrated with SC for end-to-end (E2E) training to resist wireless noise while improving the overall performance of SC [16, 10, 17, 18].
@@
Unlike SC for text or images, video-based SC presents greater challenges due to the extra temporal correlations presented in videos. Building on traditional video compression techniques, earlier studies [19, 11, 20, 21] break down video transmission into the sequential transmission of several frames using conditional coding. The current frame is modeled as the conditional distribution with respect to the adjacent reference frames. Then, frames are encoded and decoded sequentially based on the reference frames in actual transmission. Some studies [22, 23] segment frames into backgrounds and key points/segments for semantic extraction and transmission. However, these approaches lack spatiotemporal modeling of the whole video, leading to inefficient semantic extraction. Overall, current video-based SC research mainly focuses on video reconstruction, with few investigations and developments for other intelligent functionalities. Furthermore, research into multimodal video-based SC has yet to be extensively explored.
@@
Compared with image-based visual question answering, VideoQA includes a broader range of question types. It involves not only recognition of visual objects, actions, and events, but also reasoning of spatiotemporal and causal relationships, making it more challenging [24]. The key of VideoQA is the understanding of video contents with questions, which drives extensive research on how to effectively handle videos and questions [25, 26, 27, 28, 29, 30]. Some works jointly extract frame features and motion features to get effective video representations [25, 26, 27]. The generic backbone like Vision Transformers are used to obtain general video representations in [28, 29]. Moreover, innovative loss and training methods are developed to align video and question features for multimodal fusion [26, 28, 29, 30]. However, the video features extracted by traditional VideoQA methods are usually of high dimensions, which may not meet bandwidth constraints in wireless networks. Channel fading and noise also affect the accurate transmission of video features, resulting in degradation of VideoQA performance.
@@
    How to mitigate the effects of wireless channel degradation and meet bandwidth limitations while maintaining effective VideoQA performance?
In this paper, we investigate an E2E multimodal SC system named VideoQA-SC for VideoQA tasks. The proposed VideoQA-SC mainly incorporates two customized modules to address the above challenges: a spatiotemporal video semantic encoder and a learning-based bandwidth-adaptive joint source-channel (JSC) encoder/decoder. Experimental results demonstrate that the proposed VideoQA-SC achieves noise robustness and bandwidth efficiency.
@@
    An E2E SC System for VideoQA Tasks: We propose an E2E multimodal SC system called VideoQA-SC for VideoQA tasks. VideoQA-SC exploits the efficient video semantic extraction and the bandwidth-adaptive DJSCC transmission to fully leverage video information, which is noise robustness and bandwidth efficiency with promising task performance.
@@
    Cross-Attention Based JSC Encoder/Decoder: We propose a dual-branch cross-attention Transformer structure as both the JSC encoder and decoder with the learnable rate embedding shared between the transmitter and receiver. The structure allows for progressive refinement of the semantics at both the transmitter and receiver by the cross-attention architecture.
@@
The rest of this paper is organized as follows. The system model and the process for performing VideoQA tasks are introduced in Section 1. We explain our proposed methods and the detailed network architectures in Section 2. Section 3 provides the quantified experimental results and the comparison with existing advanced methods. Finally, Section 4 summarizes this paper and gives conclusions.
Notations: In this paper, lowercase letters, e.g., ${x}$, denote scalars. Bold lowercase letters, e.g., $\mathbf{x}$, denote vectors and bold uppercase letters, e.g., $\mathbf{X}$, denote matrices or tensors. $\mathbf{I}$ denotes the identity matrix. $\mathcal{CN}(\mu,\sigma^{2})$ and $\mathcal{N}(\mu,\sigma^{2})$ denote the complex Gaussian distribution and the standard Gaussian distribution with mean $\mu$ and covariance $\sigma^{2}$, respectively. $\log_{2}$ denotes the logarithm to base $2$ and $\log$ denotes the natural logarithm. $(\cdot)^{T}$ denotes the transpose and $(\cdot)^{*}$ denotes the conjugate transpose. $\mathbb{R}$ and $\mathbb{C}$ denote the real set and the complex set, respectively. $\mathbb{E}[\cdot]$ denotes the statistical expectation operation. $\operatorname{Uniform}(a,b)$ denotes the uniform distribution with start $a$ and end $b$.
1 System Model
@@
where $a^{\star}$ is the predicted answer chosen from the candidate answers, i.e., multiple choices or a predefined global answer set, denoted as $\mathcal{A}$, and $o_{\boldsymbol{\omega}}(\cdot)$ is the VideoQA model with the learnable vector $\boldsymbol{\omega}$.
@@
1.1 Transmitter
The transmitter first extracts the low-dimensional semantics $\mathbf{Y}_{v}\in\mathbb{R}^{l_{v}\times{d}}$ from the input video $\mathbf{X}_{v}$ using the spatiotemporal semantic encoder $g(\cdot;\boldsymbol{\zeta})$ with the learnable parameter vector $\boldsymbol{\zeta}$. Then, the JSC encoder with rate predictors $f_{e}(\cdot;\boldsymbol{\theta},\boldsymbol{\epsilon})$ processes $\mathbf{Y}_{v}$ into $\mathbf{S}_{v}\in\mathbb{R}^{l_{v}\times{d}}$, part of whose channels are masked as zero. $\boldsymbol{\theta}$ and $\boldsymbol{\epsilon}$ are the learnable parameter vectors of the JSC encoder and rate predictors, respectively. Non-zero channels of $\mathbf{S}_{v}$ are flattened into continuous-valued real symbols. Finally, complex channel input symbols $\mathbf{s}_{v}^{c}\in\mathbb{C}^{n}$ are obtained by converting each two real symbols into one complex symbol. The process of $\mathbf{X}_{v}$ at the transmitter can be expressed as:
@@ 1.2 Channel
where $\mathbf{n}\in\mathbb{C}^{n}$ consists of independent and identically distributed (i.i.d.) samples that follow $\mathcal{CN}(0,\sigma^{2}_{n}\mathbf{I})$. $\sigma^{2}_{n}$ denotes the average noise power. For block fading channels, an additional channel gain $h\in\mathbb{C}^{1}$ is introduced for each $\mathbf{s}_{v}^{c}$:
$\mathbf{\hat{s}}_{v}^{c}=h\mathbf{s}_{v}^{c}+\mathbf{n}.$ (4)
1.3 Receiver
The received complex channel output symbols $\mathbf{\hat{s}}_{v}^{c}$ are first converted to real symbols, which are further unflattened and padded with zeros to form $\mathbf{\hat{S}}_{v}\in\mathbb{R}^{l_{v}\times{d}}$. The receiver then decodes the video semantics $\mathbf{\hat{Y}}_{v}\in\mathbb{R}^{l_{v}\times{d}}$ from $\mathbf{\hat{S}}_{v}$ by the JSC decoder $f_{d}(\cdot;\boldsymbol{\phi})$ with the learnable parameter vector $\boldsymbol{\phi}$. Subsequently, the receiver utilizes the multimodal fuser $u(\cdot;\boldsymbol{\nu})$ with the learnable parameter vector $\boldsymbol{\nu}$ to interact with video and text information and predicts the corresponding answer $a^{\star}$ to the question $\mathbf{X}_{q}$. The process of answer prediction at the receiver can be expressed as:
@@ 1.3 Receiver
The goal of VideoQA-SC is to maximize the average accuracy of VideoQA on testing data for a given bandwidth $B$ by optimizing all learnable parameter vectors $\boldsymbol{\zeta}$, $\boldsymbol{\theta}$, $\boldsymbol{\epsilon}$, $\boldsymbol{\phi}$ and $\boldsymbol{\nu}$, which can be formulated as:
$\max_{\boldsymbol{\zeta},\boldsymbol{\theta},\boldsymbol{\epsilon},\boldsymbol{\phi},\boldsymbol{\nu}}&\quad\mathbb{E}[\operatorname{ACC}(a,a_{\text{label}})]\\ \mathrm{s.t.}&\quad R\leq{B},\end{aligned}$ (6)
where $\operatorname{ACC}(\cdot,\cdot)$ is an indicator function (1 only if $a=a_{\text{label}}$ and 0 otherwise).
Figure 3: Temporal and spatial modeling of the $i$-th video clip. Shapes with the same color represent the same objects in different frames.
2 The Proposed Method
In this section, we describe in detail the proposed VideoQA-SC system, including the spatiotemporal semantic encoder, cross-attention based JSC encoder/decoder and learning-based adaptive bandwidth allocation. Then, we introduce the training strategy of VideoQA-SC.
2.1 Spatiotemporal Semantic Encoder
We develop a spatiotemporal semantic encoder $g(\cdot;\boldsymbol{\zeta})$ to extract the video semantics by modeling the spatial and temporal correlations of the video. The purpose of $g(\cdot;\boldsymbol{\zeta})$ is to extract the coarse-grained semantics that is beneficial to fully understand the video content. In this way, although there may be multiple receivers with different inquiries, they are able to discern the video content pertinent to their specific questions through the processed video semantics, enabling them to perform their own analysis without video recovery.

### v2

@@ Abstract
Although semantic communication (SC) has shown its potential in efficiently transmitting multimodal data such as texts, speeches and images, SC for videos has focused primarily on pixel-level reconstruction. However, these SC systems may be suboptimal for downstream intelligent tasks. Moreover, SC systems without pixel-level video reconstruction present advantages by achieving higher bandwidth efficiency and real-time performance of various intelligent tasks. The difficulty in such system design lies in the extraction of task-related compact semantic representations and their accurate delivery over noisy channels. In this paper, we propose an end-to-end SC system, named VideoQA-SC for video question answering (VideoQA) tasks. Our goal is to accomplish VideoQA tasks directly based on video semantics over noisy or fading wireless channels, bypassing the need for video reconstruction at the receiver. To this end, we develop a spatiotemporal semantic encoder for effective video semantic extraction, and a learning-based bandwidth-adaptive deep joint source-channel coding (DJSCC) scheme for efficient and robust video semantic transmission. Experiments demonstrate that VideoQA-SC outperforms traditional and advanced DJSCC-based SC systems that rely on video reconstruction at the receiver under a wide range of channel conditions and bandwidth constraints. In particular, when the signal-to-noise ratio is low, VideoQA-SC can improve the answer accuracy by 5.17% while saving almost 99.5% of the bandwidth at the same time, compared with the advanced DJSCC-based SC system. Our results show the great potential of SC system design for video applications.
Index Terms:
@@
Semantic communication, video question answering, DJSCC, bandwidth allocation, multimodal task.
I Introduction
@@
With the development of artificial intelligence (AI) technology, many edge machines deploy AI models to process information for intelligent tasks[1, 2], which support AI-empowered applications such as remote healthcare, autonomous driving, and the Internet of Things (IoT). Semantic communication (SC) is an emerging paradigm which aims to extract the semantic information in the source data and accomplish accurate meaning delivery, ultimately completing the intelligent tasks[3]. Thanks to the effective semantic extraction by deep neural networks, SC can achieve higher data compression ratios and faster execution of intelligent tasks compared to the traditional communication. As a result, SC is widely used in many intelligent applications that require low latency and high accuracy under limited bandwidth resources, e.g., IoT networks[4, 5], intelligently connected vehicle networks[6, 7] and smart factories [8].
@@
In a typical SC system, the transceiver is designed as a semantic codec (semantic encoder/decoder) represented by a neural network [9, 10, 11]. The semantic encoder at the transmitter needs to remove data redundancy and extract compact semantic representations based on the structural characteristics of the source data. The semantic decoder at the receiver aims to process received semantic information to obtain results according to the specific intelligent task.
@@
For different source data modalities (speeches, texts, images, etc.), appropriate neural network architectures are essential for semantic codecs to achieve efficient SC. Long Short-Term Memorys and Transformers can be utilized to model the sequential information for texts[12, 13], and convolutional neural networks (CNNs) can be utilized to extract local information for speeches[9, 14, 15] and images[16, 10, 17]. As a novel model architecture, diffusion-based models are also increasingly applied for SC[18, 19, 20, 21], which can learn the original data distribution through a gradual denoising process. Diffusion-based models can serve as a plug-in module to remove channel noise[18, 19] or a semantic decoder to refine the degraded signals[20, 21] at the receiver. Furthermore, deep joint source-channel coding (DJSCC) can be integrated with SC for end-to-end (E2E) training to resist wireless noise while improving the overall performance of SC[22, 24, 23]. The introduction of the encryption module can also improves the security of SC systems[25].
@@
Unlike SC for texts or images, video-based SC faces greater challenges due to the extra temporal correlations presented in videos. Building on traditional video compression techniques, earlier studies [26, 11] break down video transmission into the sequential transmission of several frames using conditional coding. The current frame is modeled as the conditional distribution with respect to the adjacent reference frames. Then, frames are encoded and decoded sequentially based on the reference frames in actual transmission. Some studies [27, 28] segment frames into backgrounds and key points/segments for semantic extraction and transmission. However, these approaches still focus on pixel-level video reconstruction without deeper exploration of information in videos.
For task-oriented video transmission, it is necessary to extract deeper semantics containing the meaning of video contents. Shao et al. [29] proposed a video transmission system TOCOM-TEM for edge video analytics tasks. The system includes task-relevant feature extraction, feature encoding and joint inference modules. To reduce the communication overhead, deterministic information bottleneck and temporal entropy models are leveraged. Nevertheless, its semantic encoding of video is still decomposed into conditional coding of sequential frames. This method lack global spatiotemporal modeling of videos, leading to inefficient semantic extraction. Wan et al. [30] proposed to use 3D-CNNs to extract spatiotemporal correlations for human activity recognition tasks, with limitations such as high computational overheads and low operational efficiency. Both these methods directly extract high-level semantics from pixel-level videos, which is challenging and inefficient. Overall, current video-based SC research mainly focuses on video reconstruction, with few investigations and developments for other intelligent functionalities. Few existing task-oriented video modeling methods are not effective for extracting high-level video semantics.
@@
Compared with image-based visual question answering, VideoQA includes a broader range of question types. It involves not only recognition of visual objects, actions, and events, but also reasoning of spatiotemporal and causal relationships, making it more challenging [31]. The key of VideoQA is to understand video contents with questions, which can be decoupled into text, video analysis and fusion. On the one hand, the multi-task (MT) learning approaches for SC can provide a useful view[32, 33, 34], which achieve efficient communication for multiple semantic tasks with multimodal data. Feature parsing and decoding with multi-exit task heads are exploited in [32] and [33] for handling various semantic tasks given the same features, respectively. Zhang et al.[34] proposed a unified MT system for SC, which can handles semantic tasks involving texts and images. In general, these approaches lack effective exploration and design for videos. On the other hand, traditional VideoQA works give some methods for video processing[35, 36, 37, 38, 39]. Some works jointly extract frame features and motion features to get effective video representations [35, 36, 37]. The generic backbone like Vision Transformers are used to obtain general video representations in [38, 39]. However, the video features extracted by traditional VideoQA methods are usually of high dimensions, which may not meet bandwidth constraints in wireless networks. Moreover, channel fading and noise also affect the accurate transmission of video features, resulting in degradation of VideoQA performance.
@@
    How to mitigate the effects of wireless channel degradation and meet bandwidth constraints while maintaining the performance of VideoQA tasks?
In this paper, we investigate an E2E SC system named VideoQA-SC for VideoQA tasks. The proposed VideoQA-SC mainly incorporates two customized modules to address the above challenges: a spatiotemporal video semantic encoder and a learning-based bandwidth-adaptive joint source-channel (JSC) encoder/decoder. Experimental results demonstrate that the proposed VideoQA-SC achieves noise robustness and bandwidth efficiency.
@@
    An E2E SC System for VideoQA Tasks: We propose an E2E SC system called VideoQA-SC for VideoQA tasks. VideoQA-SC exploits the efficient video semantic extraction and the bandwidth-adaptive DJSCC transmission to fully leverage video information, which is noise robustness and bandwidth efficiency with promising task performance.
@@
    Cross-Attention Based JSC Encoder/Decoder: We propose a dual-branch cross-attention Transformer architecture as both the JSC encoder and decoder with the learnable rate embedding shared between the transmitter and receiver. The structure allows for progressive refinement of the semantics at both the transmitter and receiver by the cross-attention mechanism.
@@
The rest of this paper is organized as follows. The system model and the process for performing VideoQA tasks are introduced in Section II. We explain our proposed methods and the detailed network architectures in Section III. Section IV provides the quantified experimental results and the comparison with existing advanced methods. Finally, Section V summarizes this paper and gives conclusions.
Notations: In this paper, lowercase letters, e.g., ${x}$, denote scalars. Bold lowercase letters, e.g., $\mathbf{x}$, denote vectors and bold uppercase letters, e.g., $\mathbf{X}$, denote matrices or tensors. $\mathbf{I}$ denotes the identity matrix. $\mathcal{CN}(\mu,\sigma^{2})$ and $\mathcal{N}(\mu,\sigma^{2})$ denote the complex Gaussian distribution and the real Gaussian distribution with mean $\mu$ and covariance $\sigma^{2}$, respectively. $\log_{2}$ denotes the logarithm to base $2$ and $\log$ denotes the natural logarithm. $(\cdot)^{T}$ denotes the transpose and $(\cdot)^{*}$ denotes the conjugate transpose. $\mathbb{R}$ and $\mathbb{C}$ denote the real set and the complex set, respectively. $\mathbb{E}[\cdot]$ denotes the statistical expectation operation. $\operatorname{Uniform}(a,b)$ denotes the uniform distribution with start $a$ and end $b$.
II System Model
@@
where $a^{\star}$ is the predicted answer chosen from the candidate answers, i.e., multiple choices or a predefined global answer set, denoted as $\mathcal{A}$, and $o_{\boldsymbol{\omega}}(\cdot)$ is the VideoQA model with the learnable parametric vector $\boldsymbol{\omega}$.
@@
II-1 Transmitter
The transmitter first extracts the low-dimensional semantics $\mathbf{Y}_{v}\in\mathbb{R}^{l_{v}\times{d}}$ from the input video $\mathbf{X}_{v}$ using the spatiotemporal semantic encoder $g(\cdot;\boldsymbol{\zeta})$ with the learnable parametric vector $\boldsymbol{\zeta}$. Then, the JSC encoder with rate predictors $f_{e}(\cdot;\boldsymbol{\theta},\boldsymbol{\epsilon})$ processes $\mathbf{Y}_{v}$ into $\mathbf{S}_{v}\in\mathbb{R}^{l_{v}\times{d}}$, part of whose channels are masked as zero. $\boldsymbol{\theta}$ and $\boldsymbol{\epsilon}$ are the learnable parametric vectors of the JSC encoder and rate predictors, respectively. Non-zero channels of $\mathbf{S}_{v}$ are flattened into continuous-valued real symbols. Finally, complex channel input symbols $\mathbf{s}_{v}^{c}\in\mathbb{C}^{n}$ are obtained by converting each two real symbols into one complex symbol. The process of $\mathbf{X}_{v}$ at the transmitter can be expressed as:
@@ II-2 Channel
where $\mathbf{n}\in\mathbb{C}^{n}$ consists of independent and identically distributed (i.i.d.) samples following $\mathcal{CN}(0,\sigma^{2}_{n}\mathbf{I})$. $\sigma^{2}_{n}$ denotes the average noise power. For Rayleigh block fading channels, an additional channel gain $h\in\mathbb{C}^{1}$ is introduced for each $\mathbf{s}_{v}^{c}$:
$\mathbf{\hat{s}}_{v}^{c}=h\mathbf{s}_{v}^{c}+\mathbf{n},$ (4)
where $h$ is sampled from the Rayleigh distribution $\operatorname{Rayleigh}(\sigma_{h})$ with fading parameter $\sigma_{h}$. Here, the statistical SNR and the current SNR for each $\mathbf{s}_{v}^{c}$ are given by $\frac{h^{2}}{\sigma_{n}^{2}}$ and $\frac{\sigma_{h}^{2}}{\sigma_{n}^{2}}$, respectively.
II-3 Receiver
The received complex channel output symbols $\mathbf{\hat{s}}_{v}^{c}$ are first converted to real symbols, which are further unflattened and padded with zeros to form $\mathbf{\hat{S}}_{v}\in\mathbb{R}^{l_{v}\times{d}}$. The receiver then decodes the video semantics $\mathbf{\hat{Y}}_{v}\in\mathbb{R}^{l_{v}\times{d}}$ from $\mathbf{\hat{S}}_{v}$ by the JSC decoder $f_{d}(\cdot;\boldsymbol{\phi})$ with the learnable parametric vector $\boldsymbol{\phi}$. Subsequently, the receiver utilizes the multimodal fuser $u(\cdot;\boldsymbol{\nu})$ with the learnable parametric vector $\boldsymbol{\nu}$ to interact with video and text information and predicts the corresponding answer $a^{\star}$ to the question $\mathbf{X}_{q}$. The process of answer prediction at the receiver can be expressed as:
@@ II-3 Receiver
The goal of VideoQA-SC is to maximize the average accuracy of VideoQA on testing data for a given bandwidth $B$ by optimizing all learnable parametric vectors $\boldsymbol{\zeta}$, $\boldsymbol{\theta}$, $\boldsymbol{\epsilon}$, $\boldsymbol{\phi}$ and $\boldsymbol{\nu}$, which can be formulated as:
$\max_{\boldsymbol{\zeta},\boldsymbol{\theta},\boldsymbol{\epsilon},\boldsymbol{\phi},\boldsymbol{\nu}}&\quad\mathbb{E}[\operatorname{ACC}(a^{\star},a_{\text{label}})]\\ \mathrm{s.t.}&\quad R\leq{B},\end{aligned}$ (6)
where $\operatorname{ACC}(\cdot,\cdot)$ is an indicator function (1 only if $a^{\star}=a_{\text{label}}$ and 0 otherwise).
III The Proposed Method
In this section, we describe the proposed VideoQA-SC system in detail, which includes the spatiotemporal semantic encoder, cross-attention based JSC encoder/decoder and learning-based adaptive bandwidth allocation. Then, we introduce the training strategy of VideoQA-SC.
III-A Spatiotemporal Semantic Encoder
We develop a spatiotemporal semantic encoder $g(\cdot;\boldsymbol{\zeta})$ to extract the video semantics by modeling the spatial and temporal correlations of videos. The purpose of $g(\cdot;\boldsymbol{\zeta})$ is to extract the coarse-grained semantics that is beneficial to fully understand the video content. In this way, although there may be multiple receivers with different inquiries, they are able to discern the video content pertinent to their specific questions through the processed video semantics, enabling them to perform their own analysis without video recovery.